# Sufficient digits and density estimation: A Bayesian nonparametric approach using generalized finite Pólya trees

### v1

@@ Abstract
This paper proposes a novel approach for statistical modelling of a continuous random variable $X$ on $[0,1)$, based on its digit representation $X=.X_{1}X_{2}\ldots$. In general, $X$ can be coupled with a random variable $N$ so that if a prior of $N$ is imposed, $(X_{1},\ldots,X_{N})$ becomes a sufficient statistics and .$X_{N+1}X_{N+2}\ldots$ is uniformly distributed. In line with this fact, and focusing on binary digits for simplicity, we propose a family of generalized finite Pólya trees that induces a random density for a sample, which becomes a flexible tool for density estimation. Here, the digit system may be random and learned from the data. We provide a detailed Bayesian analysis, including closed form expression for the posterior distribution which sidesteps the need of MCMC methods for posterior inference. We analyse the frequentist properties as the sample size increases, and provide sufficient conditions for consistency of the posterior distributions of the random density and $N$. We consider an extension to data spanning multiple orders of magnitude, and propose a prior distribution that encodes the so-called extended Newcomb-Benford law. Such a model shows promising results for density estimation of human-activity data. Our methodology is illustrated on several synthetic and real datasets.
@@ 1 Introduction
Consider a continuous random variable $X$ defined on the unit interval and let $X_{1},X_{2},\ldots$ be its binary digits such that $X=\sum_{n=1}^{\infty}X_{n}2^{-n}$. In practice, due to a round-off error or the finite arithmetic precision of computers, the data consist only of a finite number of digits. In fact, under very weak conditions, there is a coupling of $X$ with a non-negative integer-valued random variable $N$ such that $X_{1},\ldots,X_{N}$ become the sufficient digits in the sense of Møller [29] and they are independent of the remaining digits $X_{N+1},X_{N+2},\ldots$ which contain no information.
@@ 1.1 Overview on models and results
In this paper, we build on this coupling result from [29] and propose a novel approach to statistical modelling for a random variable based on its digit representation. As detailed in Theorem 3.2 and Remark 4, the coupling result can be presented in a very general setting with random digits defined in terms of nested partitions of the unit interval, where the intervals of each partition are not necessarily of equal length as explained in Definition 3.1 and Remark 2 below. Thinking of $N$ as an latent variable which determines the needed number of digits, and working with binary digits for simplicity, conditioned on $N=n$, we propose models where for $j=1,\ldots,n$, we let $X_{j}\,|\,X_{1},\ldots,X_{j-1}$ be Bernoulli-distributed with a beta-distributed probability parameter depending on $X_{1},\ldots,X_{j-1}$, and where $X_{n+1},X_{n+2},\ldots$ are independent of the first $n$ digits and contain no information. Our prior models can be seen as generalizations of Pólya trees (PT) [25, 26, 27] with a finite number of partition levels, and are therefore termed generalized finite Pólya trees of type 1 (for short GFPT1; we will introduce the type 2 models below). PT priors play an important role in non-parametric Bayesian inference, see e.g. [1, 6, 7, 16, 18, 23, 31, 33, 38]. A GFPT1 prior specifies the distribution of a random probability density function (PDF) used for the data model of exchangeable observations. This random PDF is a mixture over $N$ of functions on the intervals of the binary partitions.
@@
We provide a closed-form expression for the posterior distribution which allows for straightforward inference without resorting to Markov chain Monte Carlo algorithms. In fact, our models can be fitted to tens of millions of datapoints in a matter of a fraction of a second on a standard laptop. We investigate the frequentist properties of the Bayesian estimators obtained from the posterior distribution when assuming each data point was generated from a true PDF $f^{*}$: First, we consider the posterior distribution of $N$. Assuming that $n^{*}$ digits are sufficient for $f^{*}$ (in the sense of Theorem 3.2), we establish conditions under which the posterior of $N$ concentrates on $n^{*}$, assuming that we have observed either the data perfectly (i.e., without round-off errors), or that we have access only to the first $\bar{n}$ digits, which is arguably the most common situation in applied settings. In the latter case, we identify a curious situation where ignoring the round-off errors leads to $N$ diverging to infinity a posteriori. Instead, assuming that the round-off error is uniform leads to the correct asymptotic behaviour. Second, we consider the convergence of the random PDF, establishing consistency with respect to both the Kolmogorov-Smirnov and the Hellinger distance.
@@
We propose two generalizations of GFPT1 priors. For the first ones, termed generalized finite Pólya trees of type 2 (GFPT2), we let the nested binary partitions be random such that the lengths of the intervals are obtained from sequences of beta-distributed random variables as detailed later in Definition 4.5. This additional stochastic layer is essential: Under a fixed deterministic partition one may require an unbounded number of digits ($n^{\ast}=\infty$) even when the true PDF $f^{\ast}$ is merely piecewise constant, simply because its discontinuities do not coincide with the pre-specified nested partitions. By randomising the partitions, a GFPT2 prior circumvents this limitation and restores the finite digit sufficiency property for a broad class of densities. The second generalization, discussed in Section 6, proposes how to model a sample consisting of positive numbers by combining a GFPT1 prior with a prior distribution for the order of magnitude of each data point. We find such a model to be particularly useful for density estimation in problems where data span multiple orders of magnitude and obey the so-called extended Newcomb-Benford law which routinely happens in human data [5, 8, 22]. In brief, the Newcomb-Benford law asserts that small digits are more likely to occur as leading digit in human data, and we show in Section 4.3 how to incorporate this as prior information by extending certain GFPT1s to what we call multiscale Benford Pólya trees (MBPT).
@@
We mainly restrict attention to the case $10<X<$ 1, since any real number can be transformed to a number between 0 and 1 by a bijective mapping (e.g. $x\mapsto\exp(x)/(1+\exp(x))$) and since data consisting of positive numbers can be modelled by using MBPT priors. To cover data consisting of real numbers, MBPT priors may of course be extended by including a prior for the sign.
@@ 1.3 Outline
The paper is organized as follows. Section 2 specifies the notation used throughout this text. Section 3 explains the meaning of sufficient digits (as briefly discussed in Section 1.1). Models for data on the unit interval are studied in Section 4: Section 4.1 recalls the definition of Pólya tree priors and their finite versions, Section 4.2 introduces our generalized finite Pólya tree priors GFPT1 and GFPT2 (cf. Section 1.1), and Section 4.3 discusses the choice of hyperparameters for such priors and the common Pólya tree priors. Various theoretical results for Bayesian analysis based on GFPT1 and GFPT2 priors are established in Section 5: Section 5.1 shows how to easily simulate from the posterior distribution. Mainly restricting attention to GFPT1 priors and shortly commenting on GFPT2 priors, Section 5.2 deals with consistency results for the posterior distribution of $N$ and Section 5.3 with consistency results for the posterior density estimate (as briefly discussed in Section 1.1). Section 6 studies our generalization of GFPT1 priors via the extended Newcomb-Benford law to MBPT priors, cf. Section 1.1. Section 7 considers numerical simulation results: Section 7.1 shows how the choice of hyperparameters affects posterior inference when using GFPT1 priors and suggests reasonable default values. Section 7.2 compares the posterior density estimate when using PT and GFPT1 priors and assess whether the increased flexibility of GFPT2 priors yields practical gains over GFPT1 ones. Section 7.3 compares results for Bayesian density estimation when using GFPT1 and GFPT2 priors, base-2 and base-10 MBPT priors, and the standard prior given by a Dirichlet process mixture of Gaussian densities (see e.g. [14]). Finally, Section 8 summarizes our findings and points to future research.
@@ 2 Notation
Let =$\mathbb{N}=\{1,2,\ldots\}$, $\mathbb{N}_{0}=\{0,1,2,\ldots\}$, $\mathbb{R}=(-\infty,\infty)$, and $\mathbb{R}_{+}=(0,+\infty)$ be the sets of positive integers, non-negative integers, real numbers, and positive real numbers, respectively.
For $n\in\mathbb{N}$ and $(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, let $x_{1:n}=(x_{1},\ldots,x_{n})$, and for $k\in\{0,1\}$, identify $x_{1:n-1},k$ by $(x_{1},\ldots,x_{n-1},k)$ if $n>1$ and by $k$ if $n=1$. Set $x_{1:0}=\emptyset$, $\{0,1\}^{0}=\{\emptyset\}$, and $x_{\emptyset,k}=x_{k}$ for $k=0,1$. We may interpret $x_{1:n}$ as an ordered configuration of $n$ points if $n>0$ and as the empty point configuration if $n=0$ (this is similar to the terminology used in point process theory [30]).
@@
Denote $\{0,1\}^{\mathbb{N}}$ the set of all infinite sequences $(x_{1},x_{2},\ldots)$ with $x_{n}\in\{0,1\},n=1,2,\ldots$. Let $\mathcal{I}$ be the collection of all half-open intervals $[a,b)$ with $10\leq a<b<$ 1 (where $a$ is included and $b$ is excluded). For a set $A$ with either $A=\mathcal{I}$ or $A\subseteq\mathbb{R}$, for $(x_{1},x_{2},\ldots)\in\{0,1\}^{\mathbb{N}}$, and for intervals/numbers $z_{x_{1:j}}\in A$ with $j\in\mathbb{N}$, consider $z=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},\,j\in\mathbb{N})$ as an infinite sequence with entries $z_{x_{1:j}}$ which appear in accordance to the ordering of the $x_{1:j}$ with respect to $\prec$. For $n\in\mathbb{N}$, restricting $z$ to its first $2^{n}$ entries, we obtain the ordered finite sequence $z^{(n)}=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},\,j=1,\ldots,n)$. For notional convenience, let $z^{(n)}=\emptyset$ denote the empty set (sequence). For $n\in\mathbb{N}_{0}$, let $z^{(>n)}=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},j=n+1,n+2,\ldots)$. Thus, $z^{(>0)}=z$.
Denote the beta-distribution on $[0,1)$ with shape parameters $\alpha_{0}>0$ and $\alpha_{1}>0$ by $\mathrm{B}(\alpha_{0},\alpha_{1})$, and its PDF by $\mathrm{B}(\cdot\,|\,\alpha_{0},\alpha_{1})$. For later use, extend this PDF to $[0,1]$ by setting $\mathrm{B}(1\,|\,\alpha_{0},\alpha_{1})=0$ (or any other arbitrary value). For $k\in\{0,1\}$, $\alpha_{k}=0$, and $\alpha_{1-k}>0$, let $\mathrm{B}(\alpha_{0},\alpha_{1})$ be the degenerated probability distribution on $[0,1]$ which is concentrated at the point $k$ (so the value of $\alpha_{1-k}$ plays no role), and for $x\in[0,1]\setminus\{k\}$, let $\mathrm{B}(k\,|\,\alpha_{0},\alpha_{1})=1$ and $\mathrm{B}(x\,|\,\alpha_{0},\alpha_{1})=0$.
For numbers $x>0$ and $y>0$, the beta function is given by $\mathrm{b}(x,y)=\int_{0}^{1}t^{x-1}(1-t)^{y-1}\,\mathrm{d}t$, and we let $b(0,x)=b(x,0)=1$.
For numbers $a<b$, $\mathrm{Unif}[a,b)$ denotes the uniform distribution on $[a,b)$.
Finally, set $10^{0}=$ 1.
@@ 3 The sufficient digits
We call $I$ a NBP of $[0,1)$ (this abbreviation is explained in Remark 1 below). Finally, define the digits of every $x\in[0,1)$ as the unique sequence $(x_{1},x_{2},\ldots)\in\{0,1\}^{\mathbb{N}}$ such that $x\in I_{x_{1:n}}$ for $n=1,2,\ldots$.
@@
Henceforth, assume for every $x\in[0,1)$ with digits $(x_{1},x_{2},\ldots)$ that $\ell_{x_{1:n}}\to 0$ as $n\to\infty$. Hence, $x$ is determined by its digits and there is a one-to-one correspondence between $[0,1)$ and $\{0,1\}^{\mathbb{N}}$. We express this by writing $x=.x_{1}x_{2}\ldots$.
Definition 3.1 may be extended by allowing an interval $I_{x_{1:n}}$ to be empty. Such an extension is relevant for certain number systems, e.g. in relation to so-called pseudo $\beta$-expansions, cf. [21] and the references therein. It is rather straightforward to modify the ideas and results in this paper to such situations.
@@
The remainder of this paper considers the following setting. Let $X$ be a random variable with a PDF $f$ concentrated on $H\subseteq[0,1)$ where $H$ has Lebesgue measure 1, and let $X_{1},X_{2},\ldots$ be the random digits of $X=.X_{1}X_{2}\ldots$. Assume that $f$ is lower semi-continuous (LSC) on $H$. Indeed, this is a very mild condition, cf. [20].
@@
and letting $N$ be a random variable with PMF $p_{n}=\int_{H}g_{n}$ for $n\in\mathbb{N}_{0}$, then $g$ is LSC on $H$ and $X\,|\,N=n$ has PDF $f_{n}=g_{n}/p_{n}$ for $p_{n}>0$. In particular, (3) is the special case where each $g_{n}$ is a non-negative constant function on each interval $I_{x_{1:n}}$ such that $\sum_{n=0}^{\infty}\int_{0}^{1}g_{n}(x)\,\mathrm{d}x=1$.
@@
On the other hand, the remainder .$X_{N+1}X_{N+2}\ldots$ is an ancillary statistic, and it is shown in [29] (see also [19, 20]) that
.$X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:N}$. (4)

### v2

@@ Abstract
This paper proposes a novel approach for statistical modelling of a continuous random variable $X$ on $[0,1)$, based on its digit representation $X=.X_{1}X_{2}\ldots$. In general, $X$ can be coupled with a latent random variable $N$ so that $(X_{1},\ldots,X_{N})$ becomes a sufficient statistics and $X_{N+1}X_{N+2}\ldots$ is uniformly distributed. In line with this fact, and focusing on binary digits for simplicity, we propose a family of generalized finite Pólya trees that induces a random density for a sample, which becomes a flexible tool for density estimation. Here, the digit system may be random and learned from the data. We provide a detailed Bayesian analysis, including closed form expression for the posterior distribution which sidesteps the need of MCMC methods for posterior inference. We analyse the frequentist properties as the sample size increases, and provide sufficient conditions for consistency of the posterior distributions of the random density and $N$. We consider an extension to data spanning multiple orders of magnitude, and propose a prior distribution that encodes the so-called extended Newcomb-Benford law. Such a model shows promising results for density estimation of human-activity data. Our methodology is illustrated on several synthetic and real datasets.
@@ 1 Introduction
Consider a continuous random variable $X$ defined on the unit interval and let $X_{1},X_{2},\ldots$ be its binary digits such that $X=\sum_{n=1}^{\infty}X_{n}2^{-n}$. In practice, due to a round-off error or the finite arithmetic precision of computers, the data consist only of a finite number of digits. In fact, under very weak conditions, there is a coupling of $X$ with a non-negative integer-valued latent random variable $N$ such that $X_{1},\ldots,X_{N}$ become the sufficient digits, i.e., they are independent of the remaining digits $X_{N+1},X_{N+2},\ldots$ which do not depend on the model for $X$, and thus contain no information and are not relevant for statistical inference, cf. Møller (2025).
@@ 1.1 Overview on models and results
In this paper, we build on this coupling result from Møller (2025) and propose a novel approach to statistical modelling for a random variable based on its digit representation. As detailed in Theorem 3.2 and Remark 4, the coupling result can be presented in a very general setting with random digits defined in terms of nested partitions of the unit interval, where the intervals of each partition are not necessarily of equal length as explained in Definition 3.1 and Remark 2 below. Thinking of $N$ as a latent variable which determines the needed number of digits, and working with binary digits for simplicity, conditioned on $N=n$, we propose models where for $j=1,\ldots,n$, we let $X_{j}\,|\,X_{1},\ldots,X_{j-1}$ be Bernoulli-distributed with a beta-distributed probability parameter depending on $X_{1},\ldots,X_{j-1}$, and where $X_{n+1},X_{n+2},\ldots$ are independent of the first $n$ digits and contain no information. Such a construction, which arises naturally from our random-digit modelling approach, can be seen as a generalizations of Pólya trees (PT) (Lavine, 1992, 1994; Mauldin et al., 1992) with a finite number of partition levels, and are therefore termed generalized finite Pólya trees of type 1 (for short GFPT1; we will introduce the type 2 models below). PT priors play an important role in non-parametric Bayesian inference (Awaya and Ma, 2024; Berger and Guglielmi, 2001; Castillo, 2017; Giordano et al., 2025; Hanson, 2006; Holmes et al., 2015; Müller et al., 2015; Paddock et al., 2003; Wong and Ma, 2010). A GFPT1 prior specifies the distribution of a random probability density function (PDF) used for the data model of exchangeable observations. This random PDF is a mixture over $N$ of functions on the intervals of the binary partitions. In particular, it is easy to see that GFPT1 priors correspond to the finite PT priors of Hanson (2006) with an additional prior for the partition depth.
The novelty of our contribution does not lie in a generalization of finite Pólya trees, but in the random-digit framework that considers digits as the primary modelling object. By placing a prior on the number of sufficient digits $N$ we turn it from a fixed tuning parameter (as in standard finite PTs) into an identifiable latent quantity amenable to inference and uncertainty quantification. By Theorem 3.2, the trailing digits beyond $N$ are ancillary and, therefore, genuinely non-informative for inference; modelling only the informative prefix yields likelihoods and priors that correctly account for round-off and finite precision. Most importantly, working at the digit level exposes a direct link to significant-digit phenomena such as the Newcomb–Benford law (Berger and Hill, 2020; Clauset et al., 2009; Hill, 1995), which routinely arises in human data. In brief, the Newcomb-Benford law asserts that small digits are more likely to occur as leading digit in human data, and we show in Section 4.3 how to incorporate this as prior information by extending certain GFPT1s to what we call multiscale Benford Pólya trees (MBPT). We find that the MBPT priors provide superior performance for density estimation in problems where data span multiple orders of magnitude. Such a setting is notoriously difficult in Bayesian density estimation due to heavy tails and different scales in the densities (Tokdar and Cunningham, 2024).
@@
Focusing on GFPT1 priors, we provide a closed-form expression for the posterior distribution which allows for straightforward inference without resorting to Markov chain Monte Carlo algorithms. In fact, our models can be fitted to tens of millions of datapoints in a matter of a fraction of a second on a standard laptop. We investigate the frequentist properties of the Bayesian estimators obtained from the posterior distribution when assuming each data point was generated from a true PDF $f^{*}$: First, we consider the posterior distribution of $N$. Assuming that $n^{*}$ digits are sufficient for $f^{*}$ (in the sense of Theorem 3.2), we establish conditions under which the posterior of $N$ concentrates on $n^{*}$, assuming that we have observed either the data perfectly (i.e., without round-off errors), or that we have access only to the first $\bar{n}$ digits, which is arguably the most common situation in applied settings. In the latter case, we identify a curious situation where ignoring the round-off errors leads to $N$ diverging to infinity a posteriori. Instead, assuming that the round-off error is uniform leads to the correct asymptotic behaviour. Second, we consider the convergence of the the posterior distribution of the random PDF, establishing consistency with respect to both the Kolmogorov-Smirnov and the Hellinger distance.
@@
Beyond the MBPT prior discussed above, we propose a furhter generalization of GFPT1 priors, termed generalized finite Pólya trees of type 2 (GFPT2), where we let the nested binary partitions be random such that the lengths of the intervals are obtained from sequences of beta-distributed random variables as detailed later in Definition 4.6. This additional stochastic layer is essential: Under a fixed deterministic partition one may require an unbounded number of digits ($n^{\ast}=\infty$) even when the true PDF $f^{\ast}$ is merely piecewise constant, simply because its discontinuities do not coincide with the pre-specified nested partitions. By randomising the partitions, a GFPT2 prior circumvents this limitation and restores the finite digit sufficiency property for a broad class of densities.
@@
We mainly restrict attention to the case $0<X<1$, since any real number can be transformed to a number between 0 and 1 by a bijective mapping (e.g. $x\mapsto\exp(x)/(1+\exp(x))$) and since data consisting of positive numbers can be modelled by using MBPT priors. To cover data consisting of real numbers, MBPT priors may of course be extended by including a prior for the sign.
@@ 1.3 Outline
The paper is organized as follows. Section 2 specifies the notation used throughout this text. Section 3 explains the meaning of sufficient digits (as briefly discussed in Section 1.1). Models for data on the unit interval are studied in Section 4: Section 4.1 recalls the definition of Pólya tree priors and their finite versions, Section 4.2 introduces our generalized finite Pólya tree priors GFPT1 and GFPT2 (cf. Section 1.1), Section 4.3 discusses the choice of hyperparameters for such priors and the common Pólya tree priors, and Section 4.4 studies our generalization of GFPT1 priors via the extended Newcomb-Benford law to MBPT priors, cf. Section 1.1. Various theoretical results for Bayesian analysis based on GFPT1 and GFPT2 priors are established in Section 5: Section 5.1 shows how to easily simulate from the posterior distribution. Mainly restricting attention to GFPT1 priors and shortly commenting on GFPT2 and MBPT priors, Section 5.2 deals with consistency results for the posterior distribution of $N$ and Section 5.3 with consistency results for the posterior distribution of the random PDF (as briefly discussed in Section 1.1). Section 6 considers numerical simulation results: Section 6.1 shows how the choice of hyperparameters affects posterior inference when using GFPT1 priors and suggests reasonable default values. Section 6.2 compares the density estimate given by the (estimated) posterior expected random PDF when using PT, GFPT1, and GFPT2 priors as well as the Optional Pólya tree (Wong and Ma, 2010) and the APT model of Ma (2017), assessing whether the increased flexibility of GFPT2 priors yields practical gains over competitors. Section 6.3 compares results for Bayesian density estimation when using GFPT1 and GFPT2 priors, base-2 and base-10 MBPT priors, and the standard prior given by a Dirichlet process mixture of Gaussian densities (see e.g. Ghoshal and Vaart, 2017). Finally, Section 7 summarizes our findings and points to future research.
@@ 2 Notation
Let $\mathbb{N}=\{1,2,\ldots\}$, $\mathbb{N}_{0}=\{0,1,2,\ldots\}$, $\mathbb{R}=(-\infty,\infty)$, and $\mathbb{R}_{+}=(0,+\infty)$ be the sets of positive integers, non-negative integers, real numbers, and positive real numbers, respectively.
For $n\in\mathbb{N}$ and $(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, let $x_{1:n}=(x_{1},\ldots,x_{n})$, and for $k\in\{0,1\}$, identify $x_{1:n-1},k$ by $(x_{1},\ldots,x_{n-1},k)$ if $n>1$ and by $k$ if $n=1$. Set $x_{1:0}=\emptyset$, $\{0,1\}^{0}=\{\emptyset\}$, and $x_{\emptyset,k}=x_{k}$ for $k=0,1$. We may interpret $x_{1:n}$ as an ordered configuration of $n$ points if $n>0$ and as the empty point configuration if $n=0$ (this is similar to the terminology used in point process theory, see Møller and Waagepetersen, 2004).
@@
Denote $\{0,1\}^{\infty}$ the set of all infinite sequences $(x_{1},x_{2},\ldots)$ with $x_{n}\in\{0,1\},n=1,2,\ldots$. Let $\mathcal{I}$ be the collection of all half-open intervals $[a,b)$ with $0\leq a<b<1$ (where $a$ is included and $b$ is excluded). For a set $A$ with either $A=\mathcal{I}$ or $\emptyset\not=A\subseteq\mathbb{R}$, for $(x_{1},x_{2},\ldots)\in\{0,1\}^{\infty}$, and for intervals/numbers $z_{x_{1:j}}\in A$ with $j\in\mathbb{N}$, consider $z=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},\,j\in\mathbb{N})$ as an infinite sequence with entries $z_{x_{1:j}}$ which appear in accordance to the ordering of the $x_{1:j}$ with respect to $\prec$. For $n\in\mathbb{N}$, restricting $z$ to its first $2^{n}$ entries, we obtain the ordered finite sequence $z^{(n)}=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},\,j=1,\ldots,n)$. For notional convenience, let $z^{(n)}=\emptyset$ denote the empty set (sequence). For $n\in\mathbb{N}_{0}$, let $z^{(>n)}=(z_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},j=n+1,n+2,\ldots)$. Thus, $z^{(>0)}=z$.
Denote the beta-distribution on $[0,1)$ with shape parameters $\alpha_{0}>0$ and $\alpha_{1}>0$ by $\mathrm{Beta}(\alpha_{0},\alpha_{1})$, and its PDF by $\mathrm{Beta}(\cdot\,|\,\alpha_{0},\alpha_{1})$. For later use, extend this PDF to $[0,1]$ by setting $\mathrm{Beta}(1\,|\,\alpha_{0},\alpha_{1})=0$ (or any other arbitrary value). For $k\in\{0,1\}$, $\alpha_{k}=0$, and $\alpha_{1-k}>0$, let $\mathrm{Beta}(\alpha_{0},\alpha_{1})$ be the degenerated probability distribution on $[0,1]$ which is concentrated at the point $k$ (so the value of $\alpha_{1-k}$ plays no role), and for $x\in[0,1]\setminus\{k\}$, let $\mathrm{Beta}(k\,|\,\alpha_{0},\alpha_{1})=1$ and $\mathrm{Beta}(x\,|\,\alpha_{0},\alpha_{1})=0$.
Denote the beta function by $\mathrm{beta}(x,y)=\int_{0}^{1}t^{x-1}(1-t)^{y-1}\,\mathrm{d}t$ for $x>0$ and $y>0$. Let $\mathrm{beta}(0,x)=\mathrm{beta}(x,0)=1$ for $x>0$.
Denote $\mathrm{Unif}[a,b)$ the uniform distribution on $[a,b)$ when $a<b$.
Finally, set $0^{0}=1$.
@@ 3 The sufficient digits
We call $I$ a NBP of $[0,1)$ (this abbreviation is explained in Remark 1 below). Finally, define the digits of every $x\in[0,1)$ as the unique sequence $(x_{1},x_{2},\ldots)\in\{0,1\}^{\infty}$ such that $x\in I_{x_{1:n}}$ for $n=1,2,\ldots$.
@@
Henceforth, assume for every $x\in[0,1)$ with digits $(x_{1},x_{2},\ldots)$ that $\ell_{x_{1:n}}\to 0$ as $n\to\infty$. Hence, $x$ is determined by its digits and there is a one-to-one correspondence between $[0,1)$ and $\{0,1\}^{\infty}$. We express this by writing $x=.x_{1}x_{2}\ldots$.
Definition 3.1 may be extended by allowing an interval $I_{x_{1:n}}$ to be empty. Such an extension is relevant for certain number systems, e.g. in relation to so-called pseudo $\beta$-expansions, cf. Herbst et al. (2025) and the references therein. It is rather straightforward to modify the ideas and results in this paper to such situations.
@@
The remainder of this paper considers the following setting. Let $X$ be a random variable with a PDF $f$ concentrated on $H\subseteq[0,1)$ where $H$ has Lebesgue measure 1, and let $X_{1},X_{2},\ldots$ be the random digits of $X=.X_{1}X_{2}\ldots$. Assume that $f$ is lower semi-continuous (LSC) on $H$. Indeed, this is a very mild condition, cf. Herbst et al. (2023).
@@
and letting $N$ be a random variable with PMF $p_{n}=\int_{H}g_{n}$ for $n\in\mathbb{N}_{0}$, then $g$ is LSC on $H$ and X|N=nX\,|\,N=n has PDF $f_{n}=g_{n}/p_{n}$ for $p_{n}>0$. In particular, (3) is the special case where each $g_{n}$ is a non-negative constant function on each interval $I_{x_{1:n}}$ such that $\sum_{n=0}^{\infty}\int_{0}^{1}g_{n}(x)\,\mathrm{d}x=1$.
@@
On the other hand, the remainder $X_{N+1}X_{N+2}\ldots$ is an ancillary statistic, and it is shown in Møller (2025) (see also Herbst et al., 2023, 2024) that
$X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:N}$. (4)